Solve 3^(3x+4) - 30(3^x) + 1=0
Answer by mananth(16946)


3^(3x+4) −30×3^x +1=0
Rewrite the expression
81×3^3x −30×3x +1=0
substitute t=3^x

81t^ 3 −30t+1=0
Solve
Calculator


t≈−0.624609
t≈0.033434
t≈0.591174
Resubstitute t
3^x= ≈−0.624609
3^x = ≈0.033434
3^x = ≈0.591174
Find x and conclude